OpenStudy (anonymous):

Write the equation of the line that is parallel to the line 4x - 3y = -12 and passes through the point (-3, 4).

OpenStudy (jdoe0001):

say, if you solve " 4x - 3y = -12 " for "y" what do you end up with?

OpenStudy (anonymous):

y=-4/3x-4

OpenStudy (anonymous):

do you just fill in the y and x with (-3,4)?

OpenStudy (anonymous):

oh wait i got it its y=4/3x+8!